OpenStudy (anonymous):

Solve the system using a graph. {-2x = 2y -4 {2x - y = -5

OpenStudy (john_es):

May be this will help you, http://www.wolframalpha.com/input/?i=Plot%5B%7B-x+%2B2%2C2x%2B5+%7D%2C%7Bx%2C-3%2C3%7D%5D

OpenStudy (jdoe0001):

graph them both, where both lines meet, is the solution

OpenStudy (zale101):

do u know how to graph the line by using this method? x| y | | |

OpenStudy (zale101):

but first change the line equation into a slope form

OpenStudy (shamil98):

-2x - 2y=-4 2x - y = -5 -3y = -9 y = -3 2x +3 = -5 2x = -8 x = -4